Thermodynamics Tire Pressure

The pressure of an automobile tire is measured to be 190 kPa (gage) before a trip and 215 kPa (gage) after the trip. The atmospheric pressure at both locations is 95 kPa. If the temperature of the air before the trip is 25 deg C, what is the air temperature after the trip?
Answer #1

Since volume is constant

P/T = constant

P1/T1 = P2/T2

Also Pabs = Patm + Pgage

P1 = 95 + 190 = 285 kPa
P2 = 95 + 215 = 310 kPa
T1 = 273 + 25 = 298 K

=> T2 = (310 x 298) /285 = 324.14 K

= 324.14- 273 0C = 51.14 0C